OpenStudy (anonymous):

Evaluate (4x10^7)(3.6x10^-4

OpenStudy (anonymous):

\[(4*10^7*3.6*10^{-4}) = 4*3.6*10^3 = 14.4*10^3 = 14400\]

OpenStudy (anonymous):

The possible multiple choice answers are 1.44x10^4, 1.44x10^2, or 1.44x10^-27

OpenStudy (anonymous):

\[14.4*10^3 = 1.44*10^4\]

OpenStudy (anonymous):

All I did was move the decimal place over one space to the left and multiply by another 10.

OpenStudy (anonymous):

\[14400.0\] Move the decimal place over to the left 4 times. Each time you do so, you are dividing the number by 10. So we get. \[1.4400\] However, to cancel this out so that we get the original number, we multiply by 10^4, because we have divided by 10^4 \[1.44*10^4\]

OpenStudy (anonymous):

The population of camden, New Jersey has been decreasing by 0.12% a year on average. If this trend continues, and the population was 79,318 in 2006, estimate Camden's population in 2015. Can you help me solve this?

OpenStudy (anonymous):

\[79318*(1-.0012)^{2015-2006} = 79318*(1-.0012)^{9} = about 78465\]